# WaterSeeker: Efficient Detection of Watermarked Segments in Large Documents

### v2

@@ 1 Introduction
As the quality of text generated by large language models (LLMs) improves, these models address various practical challenges while also raising concerns such as misinformation (Liu et al., 2023b; Chen & Shu, 2024) and copyright infringement (Rillig et al., 2023). To tackle these issues, LLM watermarking technology has emerged, embedding specific information (watermarks) during text generation for accurate detection through specialized algorithms. Existing watermarking methods mainly focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson & Kirchner, 2022; Kuditipudi et al., 2024). However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods. For instance, some approaches categorize vocabulary into green and red lists, applying a bias towards the selection of green tokens during generation and using metrics like the z-score to measure their proportion throughout the entire text (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b). Other methods utilize pseudo-random numbers to guide token sampling, calculating correlations across the entire document during detection (Aaronson & Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024). Ultimately, the presence of a watermark is determined by whether these statistics exceed a specified threshold.
However, these full-text detection methods struggle when small amounts of watermarked text are mixed with large volumes of natural text due to dilution effects, as illustrated in Figure 1(a). To the best of our knowledge, the only existing work addressing this scenario is the WinMax (Kirchenbauer et al., 2024) algorithm, which examines all possible window sizes and traverses the full text for each size. However, this method suffers from high time complexity. Additionally, our self-constructed baseline, the Fixed-Length Sliding Window (FLSW) method, faces the issue of watermark strength dilution due to inflexible window sizes, which negatively impacts detection performance.
@@
(b) An illustration of the WaterSeeker algorithm, using “first locate, then detect” strategy to balance time complexity and detection performance.
@@
To address these issues, we propose a novel watermark detection method named WaterSeeker, which is adaptable to various watermarking techniques. This method achieves a superior balance between time complexity and detection performance when identifying watermarked segments within large documents. The core concept of WaterSeeker is “first locate, then detect”. As illustrated in Figure 1(b), it initially employs a low-complexity anomaly points extraction algorithm to identify suspected watermark regions. This process narrows the detection target from a long text down to a small segment that encompasses the ground truth segment. Next, a local traversal is performed on the localization result. Within each window, full-text watermark detection is conducted, and the highest confidence result is compared to a threshold for the final determination. Theoretical analysis indicates that this coarse-to-fine process has the potential to achieve optimal detection performance while maintaining the lowest possible complexity for solving this problem.
In the experiment, we compared the effectiveness of WaterSeeker and the baselines in detecting watermarked segments of random lengths inserted into large documents, as well as their time complexity. The results show that WaterSeeker significantly surpasses the baseline methods in effectively balancing time complexity and detection performance. Additionally, more detailed analysis includes an adaptability assessment of WaterSeeker under varying watermark strengths and segment lengths, as well as an evaluation of its two-stage utility via ablation study. Furthermore, we discuss WaterSeeker’s contribution to building more interpretable AI detection systems.
In summary, the contributions of this work are as follows:
  • We comprehensively define the new scenario of detecting watermarked segments in large documents, specifying the inputs and outputs of the algorithms designed to address this problem, as well as the evaluation metrics and the construction process for the test dataset.
  • We propose WaterSeeker, a novel and general watermark detection method that effectively identifies watermarked segments in large documents, addressing the challenges posed by dilution effects in non-watermarked text.
  • WaterSeeker achieves a superior balance between time complexity and detection effectiveness, with strong adaptability to varying watermark strengths and lengths, significantly outperforming existing baselines.
  • WaterSeeker not only provides accurate detection results but also offers precise localization outcomes, contributing to the development of more interpretable AI detection systems.
@@ 2 Related Work
Watermarking for large language models (LLMs) embeds subtle, algorithmically detectable signals in outputs to identify LLM-generated text, making it crucial for mitigating potential misuse of LLMs. Currently, mainstream watermarking methods involve intervening in the inference phase of LLMs by modifying the logits output or influencing the token sampling process to produce watermarked text (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c).
The KGW method (Kirchenbauer et al., 2023) modifies logits by categorizing the vocabulary into green and red lists, applying a bias towards green tokens during text generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. For detection, it first labels each token as red or green and calculates the z-score of the entire text to reflect the proportion of green tokens; if this proportion exceeds a specified threshold, the text is classified as watermarked. In subsequent work, numerous enhancements have been suggested to improve text quality (Hu et al., 2024; Wu et al., 2023; Takezawa et al., 2023), increase information capacity (Wang et al., 2024; Yoo et al., 2023; Fernandez et al., 2023), boost robustness against watermark removal attacks (Zhao et al., 2024; Liu et al., 2024b; Ren et al., 2024; He et al., 2024), adapt to low-entropy conditions (Lee et al., 2023; Lu et al., 2024), and facilitate public detection (Liu et al., 2024a; Fairoze et al., 2023).
@@
On the other hand, the Aar method (Aaronson & Kirchner, 2022) proposes utilizing pseudo-random sequences to guide the token sampling process. During each token sampling step, a pseudo-random vector $u\sim\text{Uniform}([0,1])^{|V|}$ is generated based on the previously sampled tokens. Assuming that the probability vector provided by the LLM is $p$, the method performs exponential sampling to select the token $i$ that maximizes $u_{i}^{1/p_{i}}$. Here, the watermark strength is reflected by a sampling temperature, which affects the strictness of the exponential sampling. For detection, it also employs global statistics: it sums the correlation values of each token with the pseudo-random vector and then performs gamma transformation to derive the detection confidence. Building on this, Kuditipudi et al. (2024) introduced the use of edit distance to calculate correlation, while Hou et al. (2024a) and Hou et al. (2024b) incorporated contextual semantic information to guide the sampling process, thereby enhancing detection robustness.
@@
Despite the high accuracy of watermarking algorithms for detecting LLM-generated text, previous studies have primarily focused on distinguishing between fully watermarked and non-watermarked text, overlooking the possibility that LLMs may only generate small segments within large documents. In such cases, watermark detection algorithms based on full-text statistics can fail due to the dilution effect. A few studies have mentioned copy-paste attack (Kirchenbauer et al., 2024; Yoo et al., 2023; Wang et al., 2024), where a certain proportion of watermarked text is mixed into non-watermarked content, resembling our scenario. Yoo et al. (2023) and Wang et al. (2024) tested the robustness of their methods against copy-paste attacks by mixing 10% to 50% watermarked text into non-watermarked text. However, since they did not specifically design detection mechanisms for this scenario, their results indicated a lack of robustness against such attack.
@@
In existing studies, WinMax (Kirchenbauer et al., 2024) is the only watermark detection algorithm specifically designed for this scenario. It examines all possible window sizes and traverses the entire text for each size to identify the maximum local score for threshold comparison. While this method shows some effectiveness, its high time complexity renders it impractical for real-world applications. Therefore, to achieve effective detection of watermarked segments inserted into large documents while maintaining controllable time complexity, we propose WaterSeeker, a novel and general method that uses “first locate then detect” strategy. We validate its superiority through comparisons with full-text detection, WinMax, and a self-constructed fixed-length sliding window detection method.
@@ 3.1 Definition
The problem of detecting watermarked segments in a long document is defined as follows: Let $N$ be a long text with a randomly inserted watermarked segment of length $L$ (where $L$ is a random variable). We denote the starting index of the watermarked segment in $N$ as $s$ and the ending index as $e$, such that: $L=e-s+1$. The objective is to determine the presence and location of a watermarked segment in $N$. This can be framed as a binary classification problem with additional localization, where the detection algorithm outputs:
@@ 3.2 Evaluation
  • There exists at least one detected segment $(s^{\prime},e^{\prime})$ such that the Intersection over Union (IoU) with the ground truth segment $(s,e)$ exceeds a specified threshold $\theta$:
    $\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$ (1)
Based on this evaluation, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score. Additionally, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
@@ 4.2 WinMax
WinMax (Kirchenbauer et al., 2024) is a watermark detection algorithm proposed on the basis of KGW to address copy-paste attack. It involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold. In practical applications, a minimum window size $W_{min}$ and a maximum window size $W_{max}$ are typically set to control the number of iterations during the traversal. The detection process can be described by the following formula:
$z_{\text{win-max}}=\max_{w\in[W_{min},W_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$ (2)
where $w$ is the length of the local window, and $|s|_{G,i}$ represents the count of green tokens within the $i$-th local window of length $w$. For other watermarking algorithms, such as Aar, a similar approach can also be implemented. The time complexity of the WinMax detection method is relatively high, costing $O((W_{max}-W_{min})\times N)$, where $N$ is the total length of the text. Due to the uncertain length of the watermarked segment inserted into the document, if there is a large difference between $W_{max}$ and $W_{min}$, the worst-case complexity can reach $O(N^{2})$.
@@ 4.3 Fix-Length Sliding Window
Fix-Length Sliding Window (FLSW) detection method is a self-constructed, straightforward approach to address the watermark detection problem in this scenario. It involves using a fixed-length window to traverse the entire text, calculating statistics within the local window. If the statistics within a local window exceed a specified threshold, the variable ‘has_watermarked’ is set to true. For localization, during the traversal, each pair of indices $(s,e)$ that meets the condition is recorded. Finally, a contiguous method with a certain tolerance is employed to provide the localization results. The pseudocode for the aforementioned baselines can be found in Appendix A.
Although the concept of FLSW is relatively simple and straightforward, the fixed-length nature makes it unable to adapt to varying watermark lengths. This limitation can result in dilution effects when the window size is either too large or too small. In Section 5.1, we will analyze this dilution effect in detail using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson & Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
@@ 5 Proposed Method: WaterSeeker
5.1 Theoretical Foundation: Gold Index is the Best
In this section, we will conduct a theoretical analysis to demonstrate that using the actual start and end indices (the gold index) for watermark detection yields the highest expected detection rate, using KGW and Aar as examples. As mentioned in Section 2, most watermarking algorithms are slight modifications of these two methods, maintaining the fundamental principles of statistical calculation. Therefore, analyzing these two approaches provides a broad applicability.
For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part. Let’s analyze the effect of window size $W$ on this statistic:
  • When $W<L$ (window size smaller than watermark length), we have $E[|s|_{G}]=W\gamma_{1}$. Therefore,
    $E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}},$ (3)
    where the expectation of z-score is proportional to $\sqrt{W}$.
  • When $W>L$ (window size larger than watermark length), we have $E[|s|_{G}]=L\gamma_{1}+(W-L)\gamma$. Therefore
    $E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}},$ (4)
    where the expectation of z-score is proportional to $1/\sqrt{W}$, decreasing as $W$ increases.
@@ 5.1 Theoretical Foundation: Gold Index is the Best
We start with the binomial distribution $B(W,\gamma)$ that describes the number of green tokens in a window of size $W$. For large sample sizes (typically $W>50$), we can apply the Central Limit Theorem to approximate this binomial distribution with a normal distribution $N(W\gamma,W\gamma(1-\gamma))$. We then standardize this to a z-score, which follows the standard normal distribution $N(0,1)$. The false positive rate alpha represents the area in the right tail of this distribution beyond $z^{*}$. Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$, where $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution. Solving for $z^{*}$, we get:
$z^{*}=\Phi^{(-1)}(1-\alpha),$ (5)
which is a constant value for different $W$.
@@
Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix B. Based on the above proof, we simulated the expected score of the statistic and the corresponding threshold as they vary with $W$ for $L=200$ using real data, as shown in Figure 2. It demonstrates that when the window size corresponds to the gold index, the highest detection rate can be achieved while maintaining an acceptable FPR. Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics representing watermark strength not reaching their peak or minimum, which in turn affects detection performance.
@@
Figure 2: Expected score of the statistics and the corresponding threshold across varying $W$.

### v3

# WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents

@@ Abstract
WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents
Leyi Pan1,     Aiwei Liu 1,     Yijian Lu1,     Zitian Gao2,     Yichen Di1,     Lijie Wen1thanks: Corresponding author,     Irwin King3,     Philip S. Yu4 1Tsinghua University    2The University of Sydney 3The Chinese University of Hong Kong    4University of Illinois at Chicago panly24@mails.tsinghua.edu.cn, liuaw20@mails.tsinghua.edu.cn, wenlj@tsinghua.edu.cn
@@ 1 Introduction
As large language models (LLMs) generate high-quality text, they address practical challenges but also raise concerns such as misinformation (Liu et al., 2023b; Chen and Shu, 2024) and copyright infringement (Rillig et al., 2023). LLM watermarking technology has emerged to tackle these issues by embedding specific information (watermarks) during text generation, allowing for accurate detection through specialized algorithms. Existing watermarking methods primarily focus on distinguishing fully watermarked text from non-watermarked content (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Lu et al., 2024; Aaronson and Kirchner, 2022; Kuditipudi et al., 2024). However, in real-world scenarios, LLMs often generate only brief segments within longer documents, making it essential to detect these watermarked sections among large amounts of non-watermarked text.
Most previous algorithms relied on computing statistics across the entire document for detection, which we refer to as full-text detection methods (detailed in Section 2). However, these full-text detection methods struggle when small watermarked sections are mixed with large volumes of natural text due to dilution effects, as illustrated in Figure 1(a). To the best of our knowledge, the only existing work addressing this scenario is the WinMax (Kirchenbauer et al., 2024) algorithm, which examines all possible window sizes and traverses the full text for each size. However, this method suffers from high time complexity. Additionally, our self-constructed baseline, the Fixed-Length Sliding Window (FLSW) method, faces the issue of watermark strength dilution due to inflexible window sizes, which lowers detection effectiveness.
@@
(b) An illustration of the WaterSeeker algorithm, using “first locate, then detect" strategy to balance time complexity and detection performance.
@@
To address these issues, we propose a novel watermark detection method called WaterSeeker. WaterSeeker follows a "first locate, then detect" approach, as shown in Figure 1(b). It initially employs a low-complexity anomaly points extraction algorithm to identify suspected watermark regions, narrowing the detection target from a long text to a small segment encompassing the ground truth segment. Next, a local traversal is performed on the localization result, conducting full-text watermark detection within each window and comparing the highest confidence result to a threshold for the final determination. Theoretical analysis suggests that this coarse-to-fine process has the potential to achieve optimal detection performance while maintaining the lowest possible complexity for solving this problem.
In the experiment, we compared the effectiveness and time complexity of WaterSeeker with baseline methods for detecting watermarked segments in large documents. WaterSeeker significantly outperformed the baselines in balancing time complexity and detection performance. We also assessed its adaptability to different watermark strengths and segment lengths, and evaluated its two-stage utility through an ablation study. In summary, the contributions of this work are as follows:
  • We comprehensively define a new scenario: detecting watermarked segments in large documents. This includes specifying algorithm inputs/outputs, evaluation metrics, and how to create test datasets.
  • We propose WaterSeeker, a general watermark detection method that effectively identifies watermarked segments in large documents, tackling the issues caused by dilution effects.
  • WaterSeeker outperforms baselines in achieving a superior balance between time complexity and detection effectiveness.
  • WaterSeeker not only provides accurate detection results but also pinpoints watermark locations precisely, helping to create more explainable AI detection systems (detailed in Appendix H).
@@ 2 Related Work
Currently, mainstream LLM watermarking methods involve modifying the inference phase by altering logits or influencing token sampling (Liu et al., 2023a; Pan et al., 2024; Liu et al., 2024c). The KGW family (Kirchenbauer et al., 2023; Zhao et al., 2024; Hu et al., 2024; Liu et al., 2024b; He et al., 2024; Lu et al., 2024; Liu et al., 2024a) categorizes vocabulary into green and red lists, biasing towards green tokens during generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. Detection involves calculating the z-score of green tokens among the entire document; exceeding a threshold indicates watermarking.
@@
On the other hand, the Aar family (Aaronson and Kirchner, 2022; Christ et al., 2024; Kuditipudi et al., 2024) uses pseudo-random sequences to guide token sampling. It generates a pseudo-random vector $u\sim\text{Uniform}([0,1])^{|V|}$ based on previous tokens and selects the token $i$ maximizing $u_{i}^{1/p_{i}}$, where $p$ is the LLM’s probability vector. Watermark strength is controlled by sampling temperature. For detection, it also employs global statistics: it sums the correlation values of each token with the pseudo-random vector and then performs gamma transformation to derive the detection confidence. Details of the KGW and Aar watermarking algorithms can be found in Appendix A.
@@
Despite the high accuracy of watermarking algorithms for detecting LLM-generated text, previous studies have primarily focused on distinguishing between fully watermarked and non-watermarked text, overlooking the possibility that LLMs may only generate small segments within large documents. In such cases, watermark detection algorithms based on full-text statistics can fail due to the dilution effect. A few studies have mentioned copy-paste attack (Kirchenbauer et al., 2024; Yoo et al., 2023; Wang et al., 2024), which involves mixing a portion of watermarked text with non-watermarked content, similar to our scenario. Yoo et al. (2023) and Wang et al. (2024) evaluated their methods’ robustness against copy-paste attacks by combining 10% to 50% watermarked text with non-watermarked text. However, as they did not develop specific detection mechanisms for this situation, their findings showed that their methods were not robust against this type of attack.
@@
In existing studies, WinMax (Kirchenbauer et al., 2024) is the only watermark detection algorithm specifically designed for this scenario. It examines all possible window sizes and traverses the entire text for each size to identify the maximum local score for threshold comparison. While this method shows some effectiveness, its high time complexity renders it impractical for real-world applications. Therefore, to achieve effective detection of watermarked segments inserted into large documents while maintaining controllable time complexity, we propose WaterSeeker, a novel and general method that uses “first locate then detect" strategy.
@@ 3.1 Definition
The problem of detecting watermarked segments in a long document is defined as follows: Let $N$ be a long text with a randomly inserted watermarked segment of length $L$. We denote the starting index of the watermarked segment in $N$ as $s$ and the ending index as $e$, such that: $L=e-s+1$. The objective is to determine the presence and location of a watermarked segment in $N$. This can be framed as a binary classification problem with additional localization, where the detection algorithm outputs:
[…]
@@ 3.2 Evaluation
$\text{IoU}=\cfrac{L_{\text{intersection}}}{L_{\text{union}}}>\theta.$
Based on this criteria, we will report the following metrics of the binary classification result: False Positive Rate (FPR), False Negative Rate (FNR), and F1 Score. Moreover, we will include the average IoU between the detected segments and ground truth segments as a supplementary metric, further demonstrating the accuracy of the localization.
@@ 4.2 WinMax
WinMax (Kirchenbauer et al., 2024) involves iterating through all possible window sizes, and for each window size, the entire text is traversed to calculate the z-score for each local window, taking the maximum z-score and comparing it against a specified threshold. In practical applications, a minimum window size $W_{min}$ and a maximum window size $W_{max}$ are typically set to control the number of iterations during the traversal. The detection process can be described by the following formula:
$z_{\text{win-max}}=\max_{w\in[W_{min},W_{max}]}\left(\max_{i}\left(\frac{|s|_{G,i}-\gamma w}{\sqrt{\gamma(1-\gamma)w}}\right)\right),$
where $w$ is the length of the local window, and $|s|_{G,i}$ represents the count of green tokens within the $i$-th local window of length $w$. The time complexity of WinMax is $O((W_{max}-W_{min})\times N)$, where $N$ is the total length of the text. Due to the uncertain length of the watermarked segment inserted into the document, if there is a large difference between $W_{max}$ and $W_{min}$, the worst-case complexity can reach $O(N^{2})$.
@@ 4.3 Fix-Length Sliding Window
Fix-Length Sliding Window (FLSW) is a self-constructed method that uses a fixed-length window to traverse the text, calculating statistics within each local window. The text is flagged as watermarked if any statistic score exceeds the threshold. For localization, the method records each pair of indices $(s,e)$ meeting the condition and uses a contiguous segment merging method with tolerance. While straightforward, FLSW struggles with varying watermark lengths due to its fixed-length nature, leading to dilution effects. Section 5.1 analyzes this effect using KGW (Kirchenbauer et al., 2023) and Aar (Aaronson and Kirchner, 2022) as examples, providing theoretical support for WaterSeeker.
@@ 5 Proposed Method: WaterSeeker
5.1 Theoretical Basis: Gold Index is the Best
This section provides the theoretical foundation of WaterSeeker, showing that using actual start and end indices (gold index) for watermark detection achieves the highest expected detection rate. We use KGW and Aar as examples, since most watermarking algorithms are slight modifications of these two methods. Thus, analyzing these approaches offers broad applicability.
For the KGW method, assuming $\gamma_{1}>\gamma$ is the proportion of green tokens in the watermarked part. Let’s analyze the effect of window size $W$ on this statistic: (1) When $W<L$ (window size smaller than watermark length):
$E[z_{W}]=\frac{W\gamma_{1}-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\sqrt{W}\cdot\frac{\gamma_{1}-\gamma}{\sqrt{\gamma(1-\gamma)}}.$
(2) When $W>L$ (window size larger than watermark length):
$E[z_{W}]=\frac{L\gamma_{1}+(W-L)\gamma-\gamma W}{\sqrt{\gamma(1-\gamma)W}}=\frac{L(\gamma_{1}-\gamma)}{\sqrt{\gamma(1-\gamma)W}}.$
@@ 5.1 Theoretical Basis: Gold Index is the Best
We start with the binomial distribution $B(W,\gamma)$ that describes the number of green tokens in a window of size $W$. For large sample sizes (typically $W>50$), we can approximate this as $|s|_{G}\sim N(W\gamma,W\gamma(1-\gamma))$. Therefore, we have $z\sim N(0,1)$. The false positive rate $\alpha$ represents the area in the right tail of this distribution beyond $z^{*}$. Mathematically, this is expressed as $\alpha=1-\Phi(z^{*})$. Solving for $z^{*}$, we get: $z^{*}=\Phi^{(-1)}(1-\alpha)$, which is a constant value for different $W$.
@@
We denote $\mu=E[\log(\frac{1}{1-u_{i}})]$ as the expected correction value for a single token. Similar to KGW, we assume $\mu_{1}>\mu$ represents the expected correlation value for the watermarked portion. Due to the complexity of the Gamma-Transformation, the detailed proof can be found in Appendix C. Based on the above proof, we simulated the expected score of the statistic and the corresponding threshold as they vary with $W$ for $L=200$ using real data, as shown in Figure 2. It demonstrates that when the window size corresponds to the gold index, the highest detection rate can be achieved while maintaining an acceptable FPR. Therefore, the watermark dilution effect refers to the situation where the window used for detection is not optimal, resulting in the computed statistics not reaching their peak or minimum, which in turn affects detection performance.
@@
Figure 2: Expected score of the statistics and the corresponding threshold across various $W$.